OpenStudy (anonymous):

The sum of 2 numbers is 17 one number is 5 more than the other what are the 2 numbers (this is for my baby sis... Plz answer)

OpenStudy (anonymous):

WILL MEDAL

OpenStudy (solomonzelman):

no need

OpenStudy (solomonzelman):

Lets call one number "x", and the other "y", okay?

OpenStudy (anonymous):

Ok

OpenStudy (solomonzelman):

`The sum of 2 numbers is 17` would translate into: x+y=17 (there is one equation for you)

OpenStudy (solomonzelman):

`one number is 5 more than the other ` would transalte into: x+5=y (there is the second equation for you)

OpenStudy (anonymous):

Okay i tried a chart but my sis threw the pencil at me

OpenStudy (solomonzelman):

So all you need to do, is to solve this system: x+y=17 x+5=y

OpenStudy (solomonzelman):

basically I am modeling your question with a system of equations. I am just retriceigning some labels/names to your two numbers, and then I translate your problem into equations.

OpenStudy (solomonzelman):

you can solve this system via substitution